Show that f:R +R, f(1) = x + 2.0 is not uniformly continuous on \mathbb{R}.
Answer #1

Any query then comment below...

Take noter sit amylas If(a) fly)) = xy + 2(2-3)) = 102-9XX+Y) 42(2-4) | 12-y) Inty+2) Simce a lo be any red number, ser El, T